# Evaluating Large Language Models on Time Series Feature Understanding: A Comprehensive Taxonomy and Benchmark

### v1

@@ Appendix E Time Series formatting
E.1 Additional results of time series formatting
@@
Appendix F Prompts
@@
Seasonality Prompts Prompt 1: Detection "Input:<time series>. Question: can you detect any cyclic or periodic patterns in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Fixed seasonal patterns: Regular, predictable patterns occurring at fixed intervals (e.g., daily, weekly, monthly). Varying amplitude: Seasonal patterns where the magnitude of the seasonal effect changes over time. Shifting seasonal pattern: When the timing of the seasonal pattern shifts over time. Multiple seasonal pattern: Presence of more than one seasonal pattern, such as daily and weekly patterns. Select one of the following answers: (a) the time series has a fixed seasonal pattern, (b) the time series has seasonal pattern with varying amplitude, (c) the time series has a shifting seasonal pattern, (d) the time series has multiple seasonal patterns. Only answer (a), (b), (c) or (d)"
@@
Anomaly Prompts "Input:<time series>. Prompt 1: Detection Question: can you detect any irregularities in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Outlier: data point that notably deviates from the overall pattern of the data. Step-spike: sudden, sustained change in the data level, followed by a return to the original baseline. Level shift: sudden and lasting change in the average value of the series.Temporal disruption: interval where data is missing or not recorded. Select one of the following answers that best describes the provided time series: (a) the time series has one or more outliers, (b) the time series has a step-spike, (c) the time series has a level shift, (d) the time series has a temporal disruption. Only answer (a), (b), (c) or (d)"
Structural Break Prompts "Input:<time series>. Prompt 1: Detection Question: can you detect any regime switches or structural breaks in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Regime Change: A shift in the time series data’s statistical properties, such as mean, variance, or auto-correlation, that persists over time. This change is often gradual and represents a new phase or ’regime’ in the data. Structural Break: An abrupt change in the time series data that leads to a new level or trend. This change is typically sudden and can be linked to specific events or shifts in the underlying process. Examine the provided time series data and select the correct option: (a) The time series data exhibits a Regime Change. (b) The time series data exhibits a Structural Break. Provide your answer as either (a) or (b)."
@@
Volatility Prompts "Input:<time series>. Prompt 1: Detection Question: can you detect any volatility in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Constant Volatility: The degree of variation in the time series remains consistent and predictable over time. Variable Volatility: The level of variation in the time series changes unpredictably over time, without a clear pattern or structure. Clustered Volatility: The time series exhibits periods where volatility is significantly higher or lower, with these periods tending to cluster together. Leverage Effect: The volatility of the time series tends to increase when the series experiences negative returns, reflecting an asymmetric response to negative versus positive market movements. Select one of the following answers: (a) The time series has constant volatility, (b) The time series has variable volatility, (c) The time series has clustered volatility, (d) The time series has leverage effect volatility Only answer (a), (b), (c) or (d)"
Appendix G Licenses

### v2

@@ Appendix F Time Series formatting
F.1 Additional results of time series formatting
@@
Appendix G Prompts
Information retrieval and arithmetic reasoning prompts – Zero-shot "Input:<time series>. Given the input time series, please answer the following questions and format your responses in a dictionary with the structure shown below: {’max_value’: {’value’:value, ’date’:date}, ’min_value’: {’value’:value, ’date’:date}, ’value_on_date <date>’: {’value’:value}}. Only provide the numerical value and/or the date as the answer for each question. Format the reply as a dictionary following the instruction."
Information retrieval and arithmetic reasoning prompts – CoT "Input:<time series>. Given the input time series, please provide concise and precise answers to the following questions and format your responses in a dictionary: {’max_value’: {’value’:value, ’date’:date}, ’min_value’: {’value’:value, ’date’:date}, ’value_on_date <date>’: {’value’:value}}. To ensure accuracy, let’s follow these steps: 1. Identify the maximum value and its date. 2. Identify the minimum value and its date. 3. Find the value on the specified date <date>. Note: Only provide the numerical value and/or the date as the answer for each question. Format the reply as a dictionary following the instruction. Let’s think step by step."
Trend Prompts – Zero-shot "Input:<time series>." Question 1: Detection "Question: can you detect a general upward or downward trend in this time series? Answer yes or no only." Question 2: Classification "Select one of the following answers: (a) the time series has a positive trend, (b) the time series has a negative trend. Provide your answer as either (a) or (b)."
Trend Prompts - CoT "Input:<time series>." Question 1: Detection "Question: Question: Can you detect a general upward or downward trend in this time series? Provide your reasoning and then answer ’Yes’ or ’No’. Let’s think step by step. First, observe the overall pattern of the data points. Do they generally increase or decrease over time? Consider the starting and ending points of the series. If the ending point is significantly higher or lower than the starting point, this might indicate a trend. Also, look at the intermediate points: do they show a consistent direction of movement, or are there major fluctuations that disrupt the trend? Now, based on these observations, determine if there is a consistent pattern indicating a trend. Finally, provide your answer as ’Yes’ or ’No’."," Question 2: Classification ""Select one of the following answers: (a) The time series has a positive trend, (b) The time series has a negative trend. Provide your answer as either (a) or (b). Let’s think step by step. First, identify the general direction of the data points. Do they appear to be moving upward or downward overall? Consider the slope of the line that could be drawn through the data points. A positive slope indicates an upward trend, while a negative slope indicates a downward trend. Check for consistency in the movement. Are most of the data points following this direction, or are there significant deviations? If the overall pattern is increasing, select (a). If it is decreasing, select (b)."
@@ Appendix G Prompts
Seasonality Prompts – Zero-shot Prompt 1: Detection "Input:<time series>. Question: can you detect any cyclic or periodic patterns in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Fixed-period: Regular, predictable seasonal patterns occurring at fixed intervals (e.g., daily, weekly, monthly). Shifting Period: Seasonal patterns where the length of the period shifts over time. Multiple seasonality: Presence of multiple overlapping seasonal patterns (e.g., both weekly and monthly seasonality) Select one of the following answers: (a) The time series has fixed-period seasonality, (b) The time series has a shift in seasonal pattern,  (c) The time series has multiple seasonal patterns. Only answer (a), (b) or (c)."
Seasonality Prompts – CoT Prompt 1: Detection "Input:<time series>. Question: Can you detect any cyclic or periodic patterns in this time series? Provide your reasoning and then answer ’Yes’ or ’No’. Let’s think step by step. First, observe the overall shape of the time series. Look for repeating patterns or cycles. Identify the peaks (high points) and troughs (low points) in the series. Are these peaks and troughs occurring at regular intervals? Measure the distance between these repeating points. If the intervals between them are consistent, it suggests a cyclic pattern. Also, consider the amplitude (height) of these peaks and troughs. Is the amplitude consistent or does it vary over time? Now, based on these observations, determine if there is a consistent cyclic or periodic pattern in the time series. Finally, provide your answer as ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Fixed-Period: Seasonality with a constant, unchanging period (e.g., monthly seasonality). Shifting Period: Seasonality where the length of the period shifts over time (e.g., a seasonal pattern that shifts slightly each year). Multiple Seasonality: Presence of multiple overlapping seasonal patterns (e.g., both weekly and monthly seasonality). Select one of the following answers: (a) The time series has a fixed-period seasonality, (b) The time series has a shifting-period seasonality,  (c) The time series has multiple seasonality. Let’s think step by step. First, identify if there is a repeating pattern at fixed intervals, which would indicate a fixed-period seasonality. If the timing of the pattern shifts, it’s a shifting-period seasonality. Finally, if there are two or more overlapping seasonal patterns, identify it as multiple seasonality. Compare the intervals and magnitudes of the peaks and troughs carefully to determine the correct pattern. Now, provide your final answer as either (a), (b), or (c)."
@@
Anomaly Prompts – Zero-shot "Input:<time series>. Prompt 1: Detection Question: can you detect any irregularities in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Spike: a sudden and brief deviation from the overall pattern of the data. Level shift: a sudden and lasting change in the average value of the series. Temporal disruption: an interval where data is missing or not recorded. Select one of the following answers that best describes the provided time series: (a) The time series has one or more spikes, (b) The time series has a level shift, (c) The time series has a temporal disruption. Only answer (a), (b), or (c)."
Anomaly Prompts – CoT "Input:<time series>. Prompt 1: Detection Question: Can you detect any irregularities in this time series? Provide your reasoning and then answer ’Yes’ or ’No’. Let’s think step by step. First, observe the overall pattern of the time series. Identify the general trend or pattern. Next, look for any points that deviate significantly from this overall pattern. These deviations could be much higher or lower than the rest of the data points. Consider the context of these deviations: are they isolated points, or do they occur in a sequence? Are there sudden jumps or drops that are not consistent with the trend? After examining these factors, determine if there are any significant irregularities. Finally, provide your answer as ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Spike: a sudden and brief deviation from the overall pattern of the data. Level shift: a sudden and lasting change in the average value of the series. Temporal disruption: an interval where data is missing or not recorded. Select one of the following answers that best describes the provided time series: (a) The time series has one or more spikes, (b) The time series has a level shift, (c) The time series has a temporal disruption. Let’s think step by step. First, identify if there are any points that stand out sharply from the rest of the data, which would indicate spikes. If there is a lasting change in the average value of the series, identify it as a level shift. If there are intervals where data appears to be missing or not recorded, classify it as a temporal disruption. Based on your observations, determine the type of irregularity present. Now, provide your final answer as either (a), (b), or (c)."
@@
Volatility Prompts – Zero-shot "Input:<time series>. Prompt 1: Detection Question: can you detect any volatility in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Constant Volatility: The degree of variation in the time series remains consistent and predictable over time. Trending Volatility: The level of variation in the time series shows a clear increasing or decreasing trend over time. Clustered Volatility: The time series exhibits periods where volatility is significantly higher or lower, with these periods tending to cluster together. Dynamic Volatility: The volatility of the time series changes over time in response to external factors (e.g., leverage effect where the volatility of the time series tends to increase when the series experiences negative returns). Select one of the following answers: (a) The time series has constant volatility, (b) The time series has trending volatility, (c) The time series has clustered volatility, (d) The time series has dynamic volatility. Only answer (a), (b), (c), or (d)."
Structural Break Prompts – Zero-shot "Input:<time series>. Prompt 1: Detection Question: can you detect any regime switches or structural breaks in this time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Regime Change: A shift in the time series data’s statistical properties, such as mean, variance, or auto-correlation, that persists over time. This change is often gradual and represents a new phase or ’regime’ in the data. Structural Break: An abrupt change in the time series data that leads to a new level or trend. This change is typically sudden and can be linked to specific events or shifts in the underlying process. Examine the provided time series data and select the correct option: (a) The time series data exhibits a Regime Change. (b) The time series data exhibits a Structural Break. Provide your answer as either (a) or (b)."
Fat tails Prompt – Zero-shot "Input:<time series>. Prompt 1: Detection Question: Considering the data provided, does the time series exhibit fat tails?  Fat tails refer to a higher likelihood of extreme values compared to a normal distribution,  indicating a higher probability of observing significant positive or negative deviations.  Only answer ’Yes’ or ’No’."
Stationarity Properties – Zero-shot "Input:<time series>. Prompt 1: Detection Question: Considering the data provided, is the time series stationary? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions of non-stationary types in time series data: (a) Trend Change: The time series exhibits a significant shift in its underlying trend, indicating a change in the mean over time. (b) Variance Change: The time series shows a change in its variability or spread. (c) Seasonality: The time series displays regular and predictable patterns that repeat over a certain period. (d) Trend and Seasonality: The time series exhibits both a significant underlying trend and seasonal patterns. This type combines elements of both trend changes and predictable seasonal fluctuations. Select one of the following answers based on your analysis of the time series: (a) The time series has a trend change, (b) The time series has a variance change, (c) The time series has seasonality, (d) The time series has both trend and seasonality. Only answer (a), (b), (c) or (d)."
Correlation – Zero-shot "Input:<time series>. Prompt 1: Detection Question: Considering the data provided, is there a correlation between the time series? Only answer ’Yes’ or ’No’" Prompt 2: Classification "Select one of the following answers: (a) The time series are positively correlated or (b) The time series are negatively correlated. Provide your answer as either (a) or (b)."
Cross-Correlation – Zero-shot "Input:<time series>. Prompt 1: Detection Question: Considering the data provided, is there a correlation (direct or lagged) between the two time series? Only answer ’Yes’ or ’No’." Prompt 2: Classification "Given the following definitions: Direct Correlation: The two time series show a direct, immediate relationship between their values,  where changes in one series directly influence the other in a straightforward manner. Direct Lagged Correlation: The two time series demonstrate a delayed relationship, where changes in one series influence the other after a certain lag period. Inverse Correlation: The two time series exhibit an inverse or negative relationship between their values, where an increase in one series typically leads to a decrease in the other, and vice versa. Inverse Lagged Correlation: The two time series show a relationship where changes in one series negatively influence the other after a certain lag period, suggesting that past increases in one series lead to future decreases in the other, and vice versa. Select one of the following answers that best describes the relationship between the two time series: (a) The two time series exhibit direct correlation, (b) The two time series exhibit direct lagged correlation, (c) The two time series exhibit inverse correlation, (d) The two time series exhibit inverse lagged correlation. Only answer (a), (b), (c), or (d)."
Appendix H Licenses
@@
Appendix I Datasheet
We provide a datasheet for evaluating large language models on time series feature understanding, following the framework in Gebru et al. (2021).
Table 15: Datasheet for Time Series Feature Understanding
Motivation
For what purpose was the dataset created? The dataset was created to evaluate the capabilities of Large Language Models (LLMs) in understanding and captioning time series data, specifically in detecting, classifying, and reasoning about various time series features.
Who created the dataset and on behalf of which entity? The dataset was created by the authors of this paper for the purposes of this research project.
Who funded the creation of the dataset? The creation of the dataset was funded by the coauthors employers.
Any other comment? The dataset is intended for evaluating the performance of LLMs on time series annotation and summarization tasks, highlighting both strengths and limitations.
Composition
What do the instances that comprise the dataset represent? Instances are synthetic time series data points, representing various time series features such as trends, seasonality, anomalies, and more.
How many instances are there in total? The dataset comprises 10 synthetic datasets with 5000 samples in the train split, 2000 samples in the validation split and 200 time series samples in the test set.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? The dataset is a curated sample representing a wide range of time series features and complexities.
What data does each instance consist of? Each instance is a time series data point with associated features, metadata, and annotations for trend, seasonality, anomalies, etc.
Is there a label or target associated with each instance? No. The dataset is primarily for evaluation of time series description and understanding tasks performed by LLMs.
Is any information missing from individual instances? No.
Are relationships between individual instances made explicit? No. Each instance is considered independently for the purpose of this benchmark.
Are there recommended data splits? Yes, the dataset includes splits for training, validation, and test to ensure consistent evaluation metrics.
Are there any errors, sources of noise, or redundancies in the dataset? We make efforts to remove errors and noise, but due to the complex nature of isolating time series features, there may be some redundancies.
Is the dataset self-contained, or does it link to or otherwise rely on external resources? The dataset is self-contained.
Does the dataset contain data that might be considered confidential? No. All data used in the dataset is synthetically generated.
Collection Process
How was the data associated with each instance acquired? The synthetic data was generated using predefined rules for each feature.
Was the data directly obtained from the individuals, or was it provided by third parties or obtained from publicly available sources? The data was synthesized using algorithmic generation methods.
Were the individuals in question notified about the data collection? Not applicable. The dataset does not contain individual personal data.
Did the individuals in question consent to the collection and use of their data? Not applicable. The dataset does not contain individual personal data.
If consent was obtained, were the consenting individuals provided with any mechanism to revoke their consent in the future or for certain uses? Not applicable. The dataset does not contain individual personal data.
Has an analysis of the potential impact of the dataset and its use on data subjects been conducted? Not applicable. The dataset does not contain individual personal data.
Preprocessing/Cleaning/Labeling
What preprocessing/cleaning was done? Synthetic data was generated with controlled features.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data? Yes, both raw and preprocessed data are saved for transparency and reproducibility.
Is the software used to preprocess/clean/label the instances available? Not at the moment, preprocessing scripts and tools might be made available in a project repository.
Uses
Has the dataset been used for any tasks already? Yes, the dataset has been used for evaluating LLMs on time series feature detection, classification, and arithmetic reasoning tasks.
Is there a repository that links to any or all papers or systems that use the dataset? Not at the moment.
What (other) tasks could the dataset be used for? The dataset could be used for further time series analysis, forecasting, anomaly detection, and other machine learning tasks involving time series data.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? The synthetic nature of some datasets might limit their applicability to real-world scenarios, but they are useful for controlled benchmarking.
Are there tasks for which the dataset should not be used? The dataset is not suitable for tasks requiring personal data or highly sensitive financial predictions without further analysis.
Distribution
Will the dataset be distributed to third parties outside of the entity on behalf of which the dataset was created? Yes, the dataset will be publicly available for research purposes.
How will the dataset be distributed? The dataset will be distributed via an online repository with appropriate licensing.
When will the dataset be distributed? The dataset will be available for distribution after the publication of the paper.
Will the dataset be distributed under a copyright or other intellectual property license, and/or under applicable terms of use? Yes.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? No.
Maintenance
Who is supporting/hosting/maintaining the dataset? The dataset is maintained by the research team and contributors.
How can the owner/curator/manager of the dataset be contacted? Contact details will be provided in the dataset repository.
Is there an erratum? Not yet, but any updates or errors will be documented in the repository.
Will the dataset be updated? Yes, future updates will be made to improve and expand the dataset.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances? Not applicable.
Will older versions of the dataset continue to be supported/hosted/maintained? Yes, previous versions will remain available for reference.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? Yes, contributions are welcomed via the dataset repository, and code for expanding the dataset will be provided upon request.
Ethical Considerations
Were any ethical review processes conducted (e.g., by an institutional review board)? No formal ethical review was conducted as the dataset does not contain sensitive personal information.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? No. The dataset contains time series data without any sensitive or potentially offensive content.
Does the dataset relate to people? No.
Does the dataset identify any subpopulations (e.g., by age, gender)? No.
Is it possible to identify individuals (i.e., one or more people) from the dataset? No.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or affiliations, health data)? No.
Are there any known risks to individuals that are represented in the dataset? No.
Does the dataset contain data that might be subject to GDPR or other data protection laws? No.